OpenStudy (anonymous):

Each year, Terrell buys a plane ticket to visit his grandmother for a week. Last year, he paid one hundred thirty-five dollars. This year, he paid one hundred eighty-nine dollars. What was the percent increase in the cost of his plane ticket? Show your work.

OpenStudy (anonymous):

@ganeshie8 @texaschic101 @tootsi123

OpenStudy (anonymous):

i think its 40 percent

OpenStudy (anonymous):

how though

OpenStudy (texaschic101):

(current value - starting value) / starting value * 100 (189 - 135) / 135 * 100 = 54/135 * 100 = 0.4 * 100 = 40% yep...thats right

OpenStudy (anonymous):

thanks :)

OpenStudy (texaschic101):

sure thing :)